# The $B^{+}\to K^{+}\nu\bar{\nu}$ decay as a search for the QCD axion

### v1

@@ Abstract
$\mathcal{B}_{a}<1.0\times 10^{-6}$ $\mathcal{B}_{a}<6.2\times 10^{-6}$ $\mathcal{B}_{a}<1.2\times 10^{-6}$
$|(F_{V})_{sb}|>8.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.6\times 10^{8}$ GeV $|(F_{V})_{sb}|>8.0\times 10^{8}$ GeV
Table 1: Upper limits on the axion BR and lower limits on the coupling-rescaled Peccei-Quinn scale at 90% C.L. for the different cases considered. All limits assume $m_{a}=0$.
The different cases described and the resulting bounds are summarized in Table 1, which assumes $m_{a}=0$. Results for a general value of the axion mass are shown in Fig. 2. This figure quantitatively conveys two important pieces of information. First, as one would expect, the bound is weakened by an order of magnitude in the $m_{\nu\bar{\nu}}$ bins where data and the SM $B^{+}\to K^{+}\nu\bar{\nu}$ prediction are in tension. These discrepant bins can actually be used to set both an upper and a lower bound on $\mathcal{B}_{a}$ [47, 31, 42], as shown in Fig. 4 in the Supplemental Sec. A.
Second, and less obviously, the bound depends only marginally on the assumption of new physics in $\mathcal{B}_{\nu\bar{\nu}}$—namely, on whether $\mu_{\nu\bar{\nu}}$ is set to unity or left floating. This is an important conclusion, making the $B^{+}\to K^{+}\nu\bar{\nu}$ decay a double probe: of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, testable by $\mu_{\nu\bar{\nu}}\neq 1$, and of new light, elusive particles, produced as $B\to Ka$, with a BR sensitivity proportional to the total $\mathcal{B}_{\nu\bar{\nu}}$ uncertainty [48]. This second point is also conveyed by the similarity between the bounds in the first and second rows of Table 1. In the case where $\mathcal{B}_{\nu\bar{\nu}}$ is fixed to its SM value, we verify that varying it by as much as 10% has a negligible impact on these results. To corroborate this conclusion even further, we evaluate the bound on $\mathcal{B}_{a}$ assuming fixed values of $\mu_{\nu\bar{\nu}}$ but different from unity. As shown in Fig. 5 of the Supplemental Sec. A, the resulting bound has only a mild dependence on the specific value of the di-neutrino signal strength—even for huge beyond-SM contributions to the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude.
@@
In conclusion, we reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of $B^{+}\to K^{+}a$, with $a$ an axion ($m_{a}\simeq 0$) or ALP ($m_{a}>0$). We derive analytically the mapping between the measured $q^{2}_{\text{rec}}$ and the true $q^{2}$, allowing the Inclusive Tag Analysis to constrain axion emission without relying on internal simulations. Using public efficiencies, we set the strongest existing limit on the Peccei–Quinn scale rescaled by the fundamental axion coupling relevant to our channel. The result is largely independent of the assumed $\mu_{\nu\bar{\nu}}$, making $B^{+}\to K^{+}\nu\bar{\nu}$ a dual probe of short-distance new physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude and of the production of an undetected axion. This approach can be generalized to $B^{+}\to K^{*+}\nu\bar{\nu}$—and to yet other channels—as soon as Belle II measurements become available.
@@ A Supplementary plots
Figure 3: Twice the negative profile log-likelihood ratio as a function of $\mu_{\nu\bar{\nu}}$ for the ITA (red), HTA (blue) or combined (black) analyses. The solid line refers to our results while the dashed line are the results from Belle II in Fig. 16 of Ref. [32]. The grey dashed lines refer to $1,2,3\sigma$ $\chi^{2}$ values with 1 degree of freedom.

### v2

@@ Abstract
$\mathcal{B}_{a}<1.0\times 10^{-6}$ $\mathcal{B}_{a}<6\times 10^{-6}$ $\mathcal{B}_{a}<1.2\times 10^{-6}$
$|(F_{V})_{sb}|>8.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>4\times 10^{8}$ GeV $|(F_{V})_{sb}|>8.0\times 10^{8}$ GeV
Table 1: Upper limits on the axion BR and lower limits on the coupling-rescaled Peccei-Quinn scale at 90% C.L. for the different cases considered. All limits assume $m_{a}=0$. See text for more details on the significant digits quoted for the different tagging strategies.
The different cases described and the resulting bounds are summarized in Table 1, which assumes $m_{a}=0$  [^7] Ref. [53] obtains a bound slightly above $1\times 10^{-6}$ for the branching ratio in the low-$m_{a}$ regime, based on a projected dataset of 0.5 ab-1. Since Ref. [53] predates Ref. [32], a detailed numerical comparison of the limits is not straightforward. We note however that we achieve a very similar sensitivity—see results in Table 1—using 0.362ab-1 of actual data, dominated by the ITA selection.. Results for a general value of the axion mass are shown in Fig. 2. This figure quantitatively conveys two important pieces of information. First, as one would expect, the bound is weakened by an order of magnitude in the $m_{\nu\bar{\nu}}$ bins where data and the SM $B^{+}\to K^{+}\nu\bar{\nu}$ prediction are in tension. These discrepant bins can actually be used to set both an upper and a lower bound on $\mathcal{B}_{a}$ [49, 31, 42, 50], as shown in Fig. 5 in Appendix A.
Second, and less obviously, the bound depends only marginally on the assumption of new physics in $\mathcal{B}_{\nu\bar{\nu}}$—namely, on whether $\mu_{\nu\bar{\nu}}$ is set to unity or left floating. This is an important conclusion, making the $B^{+}\to K^{+}\nu\bar{\nu}$ decay a double probe: of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, testable by $\mu_{\nu\bar{\nu}}\neq 1$, and of new light, elusive particles, produced as $B\to Ka$, with a BR sensitivity proportional to the total $\mathcal{B}_{\nu\bar{\nu}}$ uncertainty [51]. This second point is also conveyed by the similarity between the bounds in the first and second rows of Table 1. In the case where $\mathcal{B}_{\nu\bar{\nu}}$ is fixed to its SM value, we verify that varying it by as much as 10% has a negligible impact on these results. To corroborate this conclusion even further, we evaluate the bound on $\mathcal{B}_{a}$ assuming fixed values of $\mu_{\nu\bar{\nu}}$ but different from unity. As shown in Fig. 6 of Appendix A, the resulting bound has only a mild dependence on the specific value of the di-neutrino signal strength—even for huge beyond-SM contributions to the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude.
@@
Figure 3: QCD axion mass $m_{a}$ as a function of the effective flavor-violating coupling $(k_{V})_{sb}$, shown for two representative lower bounds on $(F_{V})_{sb}$ derived in this work. The QCD axion lies effectively at $m_{a}\simeq 0$ on collider scales, so the resulting limits correspond to vertical constraints in axion mass vs. coupling parameter space.
@@
We now close the circle with the discussion in the Introduction concerning the distinction between the theoretically motivated but experimentally challenging QCD axion, and generic ALPs, whose appeal lies mostly in their experimental accessibility. With our approach (based on a fully reproducible analysis strategy) we obtain novel bounds on the QCD axion—despite the fact that the QCD axion is generally much more challenging than ALPs at colliders, since its mass scale lies far below any collider “scales”, as clarified below.
More quantitatively, the axion mass is related to the $(F_{V})_{sb}$ scale as
$m_{a}=\frac{\sqrt{2}F_{\pi}m_{\pi}}{(F_{V})_{sb}(k_{V})_{sb}}\frac{\sqrt{m_{u}m_{d}}}{m_{u}+m_{d}}~,$ (13)
where $F_{\pi}=93\sqrt{2}~{\mathrm{M}eV}$. Our analysis yields lower bounds on $(F_{V})_{sb}$, with the weakest at about $4\times 10^{8}$ GeV and the strongest close to $9\times 10^{8}$ GeV, depending on the treatment of the di-neutrino signal strength and on the tagging strategy (hadronic or inclusive) (see Table 1). One can therefore study the corresponding range of $m_{a}$ values as a function of $|(k_{V})_{sb}|\lesssim 10$, as imposed by perturbativity. The result is shown in Fig. 3.
The figure shows that, over many orders of magnitude in the axion mass, the QCD axion is effectively located at $m_{a}\simeq 0$ on collider scales—meaning, collectively, the physical effective scales of the processes studied as well as the detection-induced resolution scales. Consequently, our limits act as essentially vertical constraints in axion parameter spaces of mass vs. coupling, constraining the effective flavor-violating coupling $(k_{V})_{sb}$ independently of $m_{a}$. This representation clarifies how our results should be interpreted within the broader axion-search landscape, which includes experiments that are sensitive to the axion mass range in our Fig. 3: while collider experiments are insensitive to the tiny QCD axion mass itself, they can nevertheless impose meaningful and novel constraints on its flavor structure.
In conclusion, we reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of the two-body decay $B^{+}\to K^{+}a$, with $a$ an axion ($m_{a}\simeq 0$) or a generic ALP ($m_{a}>0$). A central ingredient of this work is the analytic reconstruction of the mapping between the measured $q^{2}_{\mathrm{rec}}$ and the true $q^{2}$, allowing the Inclusive Tag Analysis to be fully used without relying on internal simulations. Combined with publicly available efficiencies, this yields the strongest existing limit on the Peccei–Quinn scale rescaled by the fundamental vector flavour-changing coupling relevant to our channel. A key outcome is that this bound is largely insensitive to assumptions about the di-neutrino signal strength $\mu_{\nu\bar{\nu}}$. As a result, the decay $B^{+}\to K^{+}\nu\bar{\nu}$ acts as a dual probe: it simultaneously constrains short-distance new physics in the $b\!\to\!s\nu\bar{\nu}$ amplitude and the production of light invisible states such as the axion—with the two effects remaining statistically independent to excellent accuracy. The strategy developed here can be applied directly to $B^{+}\to K^{*+}\nu\bar{\nu}$ and, more generally, to other Belle II channels as soon as the corresponding measurements become available, providing a model-independent framework for public-data reinterpretation in rare decays.
@@ Appendix A Supplementary plots
Figure 4: Twice the negative profile log-likelihood ratio as a function of $\mu_{\nu\bar{\nu}}$ for the ITA (red), HTA (blue), and combined (black) analyses. Solid lines correspond to our results, while dashed lines reproduce the Belle II results shown in Fig. 16 of Ref. [32]. The horizontal grey dashed lines indicate the $1,\,2,\,3\sigma\chi^{2}$ thresholds for one degree of freedom. (Left) The number of background events in each bin is fixed to its expected value $\langle n^{i}_{{\mathrm{b}kg}}\rangle$. (Right) The background normalization is floated as described in the text.